# Scalable multiphoton generation from cavity-synchronized single-photon sources

### v2

@@ I Introduction and main results
propagating along $N$ or more channels. The generation of this state with high fidelity and efficiency demands the use of nearly deterministic, nearly identical and perfectly synchronized single-photon sources (SPSs), each of them producing just one photon.
The best on-demand SPSs for this task rely on few-level quantum systems Fischer et al. (2016), which can be deterministically excited He and Barkai (2006); He (2006) and decay spontaneously, producing individual photons that are collected into the desired channels. The great experimental progress in controlling single quantum systems with cavity-enhanced light-matter interactions has allowed many realizations of nearly deterministic SPSs. A long list of setups includes single atoms McKeever et al. (2004); Hijlkema et al. (2007), single molecules Brunel et al. (1999); Ahtee et al. (2009); Rezai et al. (2018), trapped ions Almendros et al. (2009); Stute et al. (2012); Higginbottom et al. (2016); Sosnova et al. (2018), and atomic ensembles Chou et al. (2004); Matsukevich et al. (2006); Farrera et al. (2016); Ripka et al. (2018), as well as solid-state systems Aharonovich et al. (2016) such as quantum dots Santori et al. (2002); Nawrath et al. (2019); Wang et al. (2019b); Schnauber et al. (2019); Thoma et al. (2016); Dusanowski et al. (2019); Kiršanskė et al. (2017); Liu et al. (2018); Uppu et al. (2020) or color centers in diamond Rodiek et al. (2017); Sipahigil et al. (2014); Alléaume et al. (2004); Gaebel et al. (2004); Wu et al. (2007); Babinec et al. (2010); Tchernij et al. (2017); Neu et al. (2011); Schröder et al. (2011); Wang et al. (2019c). In the microwave regime, superconducting quantum circuits have been used to create nearly deterministic SPSs Houck et al. (2007); Pechal et al. (2014); Peng et al. (2016); Forn-Díaz et al. (2017); Gasparinetti et al. (2017); Pfaff et al. (2017); Eder et al. (2018); Zhou et al. (2020), which have the advantage of being externally tuneable Peng et al. (2016); Forn-Díaz et al. (2017); Pfaff et al. (2017), fast Pechal et al. (2014); Zhou et al. (2020), and readily integrated on-chip with very low losses Zhou et al. (2020). Currently, quantum dots in micro-pillar cavities are proving the best overall numbers including generation efficiencies, distinguishability, and single-photon purity Reimer and Cher (2019); Wang et al. (2019c); Uppu et al. (2020), but it is hard to manufacture many of them identically, and tuning them also remains elusive for more than two emitters Patel et al. (2010); Flagg et al. (2010); Moczała-Dusanowska et al. (2019); Ellis et al. (2018); Kambs and Becher (2018).
@@
$P_{N}=({\cal P}_{1})^{N}S_{N},$ (2)
of the probability ${\cal P}_{1}$ of generating an individual photon on each of the $N$ channels independently, times the correlation error $S_{N}$ introduced by the circuit. An ideal scheme should scale exponentially $(S_{N}=1)$, creating $N$ synchronous and completely uncorrelated photons. However, typical active demultiplexing schemes introduce errors that scale as $S_{N}=\frac{1}{N}\left[\eta^{N}+(N-1)\left(\frac{1-\eta}{N-1}\right)^{N}\right]$ Lenzini et al. (2017), where $\eta$ characterizes the switching efficiency. Imperfect switching is a critical issue in any large photonic circuit Li et al. (2015), but even in the limit of lossless switches ($\eta=1$), the demultiplexing scheme introduces a detrimental factor $S_{N}=1/N$ to the scaling in Eq. (2). This can significantly limit the achievable $N$-photon efficiencies in practical applications such as boson sampling.
Figure 1: Cavity-QED scheme for deterministically generating a large number of synchronized and independent single-photons. (a) Idealized implementation with NN two-level emitters coupled to a common driven cavity mode. Each emitter additionally couples to independent output channels, which collect the $N$ emitted single-photons on parallel for later usage. (b) Efficient on-chip implementation using circuit-QED. Here, transmon qubits realize the NN two-level emitters, each of them capacitively coupled to an independent transmission-line antenna and to a common transmission-line resonator. Controlling the shape $\Omega(t)$ of the resonator drive, we can trigger all SPSs simultaneously and on-demand. The NN emitted microwave single-photons are collected by the on-chip antennas with high efficiency. These antennas are also used to send a DC current to each qubit and calibrate their frequencies into resonance with the resonator drive.
@@
$S_{N}=1+\epsilon N(N-1).$ (3)
@@
The correction factor $\epsilon$ stems from residual cavity-induced correlations. For realistic experimental conditions at the optimal working point, we find that $\epsilon\sim 10^{-8}-10^{-6}$, making the multiphoton generation scheme scalable up to $N\sim 100-1000\ll\epsilon^{-1/2}$, depending on the implementation and the noise sources.
@@
To show the favorable scaling of the scheme in a realistic setup, we study a circuit-QED implementation with flux-tunable transmon emitters and microwave transmission lines DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014), considering dephasing noise, internal loss, and disorder. As shown in Fig. 1(b), this implementation is fully integrated on-chip, which allows the output antennas to have collection and transmission efficiencies above $99\%$ Zhou et al. (2020). For a strongly driven cavity, we show it is feasible to build microwave SPSs with single-photon purity and indistinguishability above $99\%$, as measured via standard Hanbury Brown and Twiss (HBT), and Hong-Ou-Mandel (HOM) experiments. Most importantly, we predict an overall single-photon efficiency or brightness of ${\cal P}_{1}\gtrsim 0.99$, a synchronized $30$-photon generation efficiency of $P_{30}\sim({\cal P}_{1})^{30}\sim 0.72$, and even $P_{100}\sim 0.33$ for $N=100$ SPSs. This means that for pulsed microwave SPSs with a repetition rate of $R\sim 0.4$ MHz, we can achieve $30$-photon generation rates of $C_{30}\sim 300$ kHz, and $100$-photon rates of $C_{100}\sim 100$ kHz. Remarkably, this is more than seven orders of magnitude more efficient than state-of-the-art boson sampling experiments with up to $N=14$ photons Wang et al. (2019a). The performance of our synchronized multiphoton generation scheme benefits from the high efficiency of the superconducting circuit implementation, as well as from the absence of switches and other optical elements. We also show that our predictions are robust to small inhomogeneities and disorder in the SPS parameters.
@@
The paper is organized as follows. In Sec. II we introduce the setup, model, main approximations, and identify the optimal parameter regime of the scheme. In Sec. III, we introduce realistic parameter sets for a circuit-QED implementation and quantify its performance via the single- and multiphoton generation efficiencies, single-photon purity, and indistinguishability. The most important results of the paper are presented in Sec. IV, where we characterize the scalability of the synchronized $N$-photon generation. In Sec. V we analyze the effect of inhomogeneities and disorder in the SPS parameters, and finally in Sec. VI we summarize our conclusions. We complement the discussion by analytical and numerical methods to quantify the photon counting and correlations, which are briefly presented in Appendices A, B, and C.
@@ II.1 General cavity-QED model
Additionally, each qubit $j$ is coupled to an independent decay channel or “antenna”, which collects the emitted photons [cf. Fig.1(a) and (b)]. We describe this qubit-antenna interaction in the the Born-Markov approximation, introducing the $\gamma_{j}$ at which a qubit deposits photons into its own antenna. We also introduce a photon loss rate $\gamma_{\rm loss}^{j}$ characterizing the emission of photons into any other unwanted channel. Finally, we consider the cavity decay rate $\kappa$ and introduce white noise dephasing rates $\gamma_{\phi}^{j}$ on each of the qubits. The complete dynamics of this open system is described with a master equation,
@@
The idealized setup in Fig. 1(a) admits various implementations. The two-level systems $\sigma_{j}^{-}$ could be neutral atoms Welte et al. (2018) or ions Casabone et al. (2015) trapped inside an optical cavity field $a$ that is localized between macroscopic mirrors. In this prototypical cavity-QED implementation the individual decay channels of each qubit would require the use of high-aperture lenses Araneda et al. (2018) or tapered nano-fibers Reitz et al. (2013) to collect the photons independently, which is very challenging to integrate and to scale-up with high collection efficiencies. Nanophotonic structures such as photonic crystals Lodahl et al. (2015) or integrated photonic circuits Antón et al. (2019); Mehta et al. (2020) are other promising platforms to realize our setup. The common mode and the independent output channels can be integrated and scaled up. In this scenario, the main limitation arise from the creation of nearly identical or tuneable emitters such as quantum dots Ellis et al. (2018), or alternatively the trapping of many atoms Goban et al. (2015) or ions Mehta et al. (2020) near these surfaces.
However, in this work we will focus on circuit-QED to discuss an efficient and scalable integration of multiple two-level emitters with individual decay channels. As sketched in Fig. 1(b), one may use superconducting transmon qubits as quantum emitters DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014). These qubits may be capacitively coupled to both a common transmission-line resonator, as well as to individual transmission-line waveguides that efficiently collect the microwave single-photons Zhou et al. (2020). In Secs. III and IV we analyze in detail the performance and scalability of this superconducting circuit platform.
@@ II.2 Cavity-mediated synchronization and residual correlations
The key mechanism to achieve an efficient cavity-mediated synchronization of the SPSs is to generate a large coherent state $|\alpha|\gg 1$ in the cavity mode. This can be achieved via a very strong drive, $|\Omega_{0}|\gg\kappa,|\Delta|$, where $\Delta=\omega_{d}-\omega_{c}$ is the cavity-drive detuning. We model the resulting state using a displacement of the Fock operator
@@
As shown in App. A.1, our control of the cavity displacement translates into a cavity-mediated driving on all coupled qubits, given by the effective Hamiltonian,
@@
This is the mechanism that allows us to excite all SPSs simultaneously. In order for our cavity mediated control to succeed, the qubit back-action, cavity-induced interactions, and correlations between qubits must be suppressed. To do so, we restrict the couplings and detunings to keep the system in the bad-cavity limit
@@
where $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ is the detuning between qubit $j$ and cavity drive. These conditions ensure that the cavity reaches a steady state (9), where small quantum fluctuations $\delta a\ll|\alpha_{\rm ss}|$ can be adiabatically eliminated and fast oscillations neglected. As shown in App. A.2, the dynamics of the qubit’s reduced state $\tilde{\rho}(t)={\rm Tr}_{c}\{\rho(t)\}$ can then be modeled by an effective master equation, which in the rotating frame with the drive frequency $\omega_{d}$ reads
$\dot{\tilde{\rho}}(t)={}$ $-i[\tilde{H}(t),\tilde{\rho}]+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+{\cal{D}}\left[\sum_{j=1}^{N}\sqrt{\gamma^{j}_{\rm cm}}\sigma_{j}^{-}\right]\tilde{\rho}$
$+\sum_{j=1}^{N}\gamma_{\rm loss}^{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\tilde{\rho}$, (12)
with the effective Hamiltonian $\tilde{H}(t)$ given by
@@
Here, the cavity-mediated driving $\Omega_{\rm cm}^{j}$ reads
@@
with the steady state coherent amplitude given in Eq. (9). In addition, the qubits experience a Lamb-shift
@@
The cavity also induces a collective decay or “superradiance” on the qubits, described by a collective jump operator $W_{\rm cm}=\sum_{j=1}^{N}\sqrt{\gamma^{j}_{\rm cm}}\sigma^{-}_{j}.$ The cavity-mediated decay rates
$\gamma^{j}_{\rm cm}={}$ $\frac{(g_{j})^{2}\kappa}{(\kappa/2)^{2}+\Delta^{2}},$ (17)
combined with the local decay and dephasing rates $\gamma_{j}$, $\gamma_{j}^{\rm loss}$, and $\gamma_{\phi},$ describe the incoherent dynamics of the qubit, as shown in Eq. (12). We will see that, for an optimal performance of the synchronized $N$-photon emitter, the residual cavity-mediated interactions $J_{\rm cm}^{jl}$ and decay $\gamma_{\rm cm}^{j}$ must be much smaller than the cavity-mediated driving $\Omega_{\rm cm}^{j}.$

### v3

@@ I Introduction and main results
propagating along $N$ or more channels. The generation of this state with high fidelity and efficiency demands the use of nearly deterministic, nearly identical, and perfectly synchronized single-photon sources (SPSs), each of them producing just one photon.
The best on-demand SPSs for this task rely on few-level quantum systems Fischer et al. (2016), which can be deterministically excited He and Barkai (2006a, b) and decay spontaneously, producing individual photons that are collected into the desired channels. The great experimental progress in controlling single quantum systems with cavity-enhanced light-matter interactions has allowed many realizations of nearly deterministic SPSs. A long list of setups includes single atoms McKeever et al. (2004); Hijlkema et al. (2007), single molecules Brunel et al. (1999); Ahtee et al. (2009); Rezai et al. (2018), trapped ions Almendros et al. (2009); Stute et al. (2012); Higginbottom et al. (2016); Sosnova et al. (2018), and atomic ensembles Chou et al. (2004); Matsukevich et al. (2006); Farrera et al. (2016); Ripka et al. (2018), as well as solid-state systems Aharonovich et al. (2016) such as quantum dots Santori et al. (2002); Nawrath et al. (2019); Wang et al. (2019b); Schnauber et al. (2019); Thoma et al. (2016); Dusanowski et al. (2019); Kirsanske et al. (2017); Liu et al. (2018); Uppu et al. (2020) or color centers in diamond Rodiek et al. (2017); Sipahigil et al. (2014); Alléaume et al. (2004); Gaebel et al. (2004); Wu et al. (2007); Babinec et al. (2010); Tchernij et al. (2017); Neu et al. (2011); Schröder et al. (2011); Wang et al. (2019c). In the microwave regime, superconducting quantum circuits have been used to create nearly deterministic SPSs Houck et al. (2007); Pechal et al. (2014); Peng et al. (2016); Forn-Díaz et al. (2017); Gasparinetti et al. (2017); Pfaff et al. (2017); Eder et al. (2018); Zhou et al. (2020), which have the advantage of being externally tuneable Peng et al. (2016); Forn-Díaz et al. (2017); Pfaff et al. (2017), fast Pechal et al. (2014); Zhou et al. (2020), and readily integrated on-chip with very low losses Zhou et al. (2020). Currently, quantum dots in micro-pillar cavities provide the best overall numbers including generation efficiencies, distinguishability, and single-photon purity Reimer and Cher (2019); Wang et al. (2019c); Uppu et al. (2020), but it is hard to manufacture many of them identically, and tuning them also remains elusive for more than two emitters Patel et al. (2010); Flagg et al. (2010); Moczała-Dusanowska et al. (2019); Ellis et al. (2018); Kambs and Becher (2018).
@@
${\color[rgb]{0,0,0}C_{N}=R}({\cal P}_{1})^{N}S_{N},$ (2)
with ${\cal P}_{1}$ the probability of generating an individual photon pulse on each of the $N$ channels independently, and $S_{N}$ the correlation error introduced by the synchronizing circuit. An ideal scheme should scale exponentially $(S_{N}=1)$, creating $N$ synchronous and completely uncorrelated photons. However, typical active demultiplexing schemes introduce errors that scale as $S_{N}=\frac{1}{N}\left[\eta^{N}+(N-1)\left(\frac{1-\eta}{N-1}\right)^{N}\right]$ Lenzini et al. (2017), where $\eta$ characterizes the switching efficiency. Imperfect switching is a critical issue in any large photonic circuit Li et al. (2015), but even in the limit of lossless switches ($\eta=1$), the demultiplexing scheme introduces a detrimental factor $S_{N}=1/N$ to the scaling in Eq. (3). This can significantly limit the achievable $N$-photon generation rates in practical applications requiring $N\gg 1$ such as single-photon boson sampling.
Figure 1: Cavity-QED scheme for deterministically generating a large number of synchronized and independent single-photons. (a) Idealized implementation with NN two-level emitters coupled to a common driven cavity mode. Each emitter additionally couples to independent output channels, which collect the $N$ emitted single photons in parallel for later usage. (b) Efficient on-chip implementation using circuit-QED. Here, transmon qubits realize the NN two-level emitters, each of them capacitively coupled to an independent transmission-line antenna and to a common transmission-line resonator. Controlling the shape $\Omega(t)$ of the resonator drive, we can trigger all SPSs simultaneously and on demand. The NN emitted microwave single-photons are collected by the on-chip antennas with high efficiency. These antennas are also used to send a DC current to each qubit and calibrate their frequencies into resonance with the resonator drive.
@@
${\color[rgb]{0,0,0}P_{N}=({\cal P}_{1})^{N}(1+D_{N})},$ (3)
except for a small error $D_{N}=[P_{N}-({\cal P}_{1})^{N}]/({\cal P}_{1})^{N}\ll 1$, characterizing the residual correlation or dependence between photons. This small deviation is shown to be quadratic in $N$:
${\color[rgb]{0,0,0}D_{N}}=\epsilon N(N-1),$ (4)
@@
with the correction factor $\epsilon$ stemming from residual cavity-induced correlations. We work out the parameter conditions to reach $\epsilon\sim 10^{-4}-10^{-7}$, making the multiphoton generation scheme scalable up to $N\sim 100-1000\ll\epsilon^{-1/2}$, depending on the implementation and the noise sources. The $N$-photon generation rate is then obtained as $C_{N}=RP_{N}$ with $R$ the repetition rate.
@@
To show the favorable scaling of the scheme in a realistic setup, we study a circuit-QED implementation with flux-tunable transmon emitters and microwave transmission lines DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014), considering dephasing noise, internal loss, and disorder. As shown in Fig. 1(b), this implementation is fully integrated on-chip, which allows the output antennas to have collection and transmission efficiencies above $99\%$ Zhou et al. (2020). Moreover, the use of the cavity as a common synchronizer has a low overhead compared to having and calibrating a large number of $N$ independent control lines and driving sources, which may lead to serious limitations in the wiring and the scalability of the superconducting chip Tamate et al. (2022). Using state-of-the-art parameters, we show it is feasible to build microwave SPSs with single-photon purity and indistinguishability of $99\%$, as measured via standard Hanbury Brown and Twiss (HBT), and Hong-Ou-Mandel (HOM) experiments. Most importantly, we predict an overall single-photon probability or brightness of ${\cal P}_{1}\gtrsim 0.99$ and the possibility to efficiently synchronize up to $N=100$ SPSs with large multi-photon probabilities such as $P_{10}\sim 0.90$, $P_{30}\sim 0.72$, and $P_{100}\sim{\color[rgb]{0,0,0}0.16}$, for $N=10$, $30$, and $100$ SPSs, respectively. This means that for pulsed microwave SPSs with a repetition rate of $R\sim 0.4$ MHz, we can achieve $30$-photon generation rates of $C_{30}\sim{\color[rgb]{0,0,0}700}$ kHz, and $100$-photon rates of $C_{100}\sim{\color[rgb]{0,0,0}200}$ kHz. Remarkably, this is more than seven orders of magnitude more efficient than state-of-the-art boson sampling experiments with up to $N=14$ single photons Wang et al. (2019a). The performance of our synchronized multiphoton generation scheme benefits from the high efficiency of the superconducting circuit implementation, as well as from the absence of switches and other optical elements. We show that our predictions are robust to small inhomogeneities in the SPS parameters, and to the multi-mode structure of the synchronizing cavity. Finally, we also consider the parameter conditions to synchronize up to $N=1000$ SPSs, which may be possible in more advanced circuit QED setups with weak and highly symmetric cavity-emitter couplings Huang et al. (2021); Hazra et al. (2021).
@@
The paper is organized as follows. In Sec. II we introduce the setup, model, main approximations, and identify the optimal parameter regime of the scheme. In Sec. III, we introduce realistic parameter sets for a circuit-QED implementation and quantify its performance via the single- and multiphoton generation efficiencies, single-photon purity, and indistinguishability. The most important results of the paper are presented in Sec. IV, where we characterize the scalability of the synchronized $N$-photon generation. In Sec. V we analyze the effects of inhomogeneities and disorder in the SPS parameters and in Sec. VI the effects of the multi-mode nature of the cavity. Finally in Sec. VII we summarize our conclusions. We complement the discussion with analytical and numerical methods to quantify the photon counting and photon correlations, which are presented in Appendices A, B, and C.
@@ II.1 General cavity-QED model
Additionally, each qubit $j$ is coupled to an independent decay channel or “antenna”, which collects the emitted photons [cf. Fig.1(a) and (b)]. We describe this qubit-antenna interaction in the Born-Markov approximation, introducing the $\gamma_{j}$ at which a qubit deposits photons into its antenna. We also introduce a photon loss rate $\gamma_{\rm loss}^{j}$ characterizing the emission of photons into any other unwanted channel. Finally, we consider the cavity decay rate $\kappa$ and introduce white noise dephasing rates $\gamma_{\phi}^{j}$ on each of the qubits. The complete dynamics of this open system are described with a master equation,
@@
The idealized setup in Fig. 1(a) admits various implementations. The two-level systems $\sigma_{j}^{-}$ could be neutral atoms Welte et al. (2018) or ions Casabone et al. (2015) trapped inside an optical cavity field $a$ that is localized between macroscopic mirrors. In this prototypical cavity-QED implementation the individual decay channels of each qubit would require the use of high-aperture lenses Araneda et al. (2018) or tapered nano-fibers Reitz et al. (2013) to collect the photons independently, which is very challenging to integrate and to scale-up with high collection efficiencies. Nanophotonic structures such as photonic crystals Lodahl et al. (2015) or integrated photonic circuits Antón et al. (2019); Mehta et al. (2020) are other promising platforms to realize our setup. The common mode and the independent output channels can be integrated and scaled up. In this scenario, the main limitation arises from the creation of nearly identical or tuneable emitters such as quantum dots Ellis et al. (2018), or the trapping of many atoms Goban et al. (2015) or ions Mehta et al. (2020) near these surfaces.
However, in this work we will focus on circuit-QED to discuss an efficient and scalable integration of multiple two-level emitters with individual decay channels. As sketched in Fig. 1(b), one may use superconducting transmon qubits as quantum emitters DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014). These qubits may be capacitively coupled to both a common transmission-line resonator, as well as to individual transmission-line waveguides that efficiently collect the microwave single-photons Zhou et al. (2020). In Secs. III and IV we analyze in detail the state-of-art parameters, the performance and the scalability of this superconducting circuit platform.
@@ II.2 Cavity-mediated synchronization and residual correlations
The key mechanism to achieve an efficient cavity-mediated synchronization of the SPSs is to generate a large coherent state $|\alpha|\gg 1$ in the cavity mode. This can be achieved via a strong drive, $|\Omega_{0}|\gg\kappa,|\Delta|$, where $\Delta=\omega_{d}-\omega_{c}$ is the cavity-drive detuning. We model the resulting state using a displacement of the Fock operator
@@
As shown in App. A.2, our control of the cavity displacement translates into a cavity-mediated driving on all coupled qubits, given by the effective Hamiltonian,
@@
This is the mechanism that allows us to excite all SPSs simultaneously. For our cavity-mediated control to succeed, the qubit back-action, cavity-induced interactions, and correlations between qubits must be suppressed. To do so, we restrict the couplings and detunings to keep the system in the bad-cavity limit
@@
where $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ is the detuning between qubit $j$ and cavity drive. These conditions ensure that the cavity reaches a steady state (10), where small quantum fluctuations $\delta a\ll|\alpha_{\rm ss}|$ can be adiabatically eliminated and fast oscillations neglected. As shown in App. A.3, the dynamics of the qubit’s reduced state $\tilde{\rho}(t)={\rm Tr}_{c}\{\rho(t)\}$ can then be modeled by an effective master equation, which in the rotating frame with the drive frequency $\omega_{d}$ reads
$\dot{\tilde{\rho}}(t)={}$ $-i[\tilde{H}(t),\tilde{\rho}]+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+{\color[rgb]{0,0,0}\sum_{j,l=1}^{N}\gamma^{jl}_{\rm cm}\tilde{{\cal D}}[\sigma_{j}^{-},\sigma_{l}^{-}]\tilde{\rho}}$
$+\sum_{j=1}^{N}\gamma_{\rm loss}^{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\tilde{\rho}$. (13)
Here, the effective Hamiltonian $\tilde{H}(t)$ has the form,
@@
with $\Omega_{\rm cm}^{j}$ the cavity-mediated driving given by
@@
with the steady state coherent amplitude $|\alpha_{\rm ss}|$ given in Eq. (10). In addition, the qubits experience a Lamb-shift
@@
The cavity also induces a collective decay or “superradiance” on the qubits, described by a generalized Lindblad term $\tilde{{\cal D}}[x,y]\rho=x\rho y^{\dagger}-(y^{\dagger}x\rho+\rho y^{\dagger}x)/2$ in Eq. (13) and cavity-mediated decay rates
${\color[rgb]{0,0,0}\gamma^{jl}_{\rm cm}}={}$ $\frac{{\color[rgb]{0,0,0}g_{j}g_{l}}\kappa}{(\kappa/2)^{2}+\Delta^{2}}.$ (18)
Finally, the effective dynamics of the qubits also include local decay and dephasing rates $\gamma_{j}$, $\gamma_{j}^{\rm loss}$, and $\gamma_{\phi}$, as described by the normal Lindblad terms in Eq. (13).